### v4

@@ Abstract
Pre-trained vision transformers have strong representation benefits to various downstream tasks. Recently, many parameter-efficient fine-tuning (PEFT) methods have been proposed, and their experiments demonstrate that tuning only 1% extra parameters could surpass full fine-tuning in low-data resource scenarios. However, these methods overlook the task-specific information when fine-tuning diverse downstream tasks. In this paper, we propose a simple yet effective method called “Salient Channel Tuning” (SCT) to leverage the task-specific information by forwarding the model with the task images to select partial channels in a feature map that enables us to tune only 1/8 channels leading to significantly lower parameter costs. Experiments outperform full fine-tuning on 18 out of 19 tasks in the VTAB-1K benchmark by adding only 0.11M parameters of the ViT-B, which is 780 fewer than its full fine-tuning counterpart. Furthermore, experiments on domain generalization and few-shot learning surpass other PEFT methods with lower parameter costs, demonstrating our proposed tuning technique’s strong capability and effectiveness in the low-data regime. The code will be available at https://github.com/showlab/SCT
@@ 1 Introduction
Large vision transformers (ViT) have achieved remarkable success in computer vision tasks (Dosovitskiy et al., 2020; Liu et al., 2021; Yuan et al., 2022; Zhou et al., 2021a; Carion et al., 2020; Li et al., 2021; Strudel et al., 2021) by using large-scale training data, such as ImageNet21K and JFT-300M, to create strong representations. However, downstream recognition tasks often lack sufficient data to train from scratch. Therefore, transferring knowledge from pre-trained ViT models can significantly reduce training difficulty and produce promising results. End-to-end full fine-tuning is commonly used to inherit these robust representations, but it faces two challenges. First, large models are prone to overfitting when tuning their massive weights on small downstream training data. Second, ViT models are too large to store all weights for each downstream task, making it infeasible to deploy fine-tuned models on resource-limited devices.
@@
To mitigate the above two challenges, tuning a subset of parameters or adopting an external trainable module to preserve the knowledge learned from pre-trained models and adapt to the downstream tasks have been proposed. As for tuning a subset of full parameters, there are two common methods: tuning the classification head (Mahajan et al., 2018; Jia et al., 2021; Chen et al., 2021b) and bias term (Cai et al., 2020). Tuning the classification head freezes the weights of the backbone network and only updates the classification head, while tuning the bias term means unfreezing the bias term in the backbone network. Both methods lead to inferior performance.
As for adopting an external trainable module, two lines of Parameter-Efficient Fine-Tuning (PEFT) methods are proposed. Firstly, adopt learnable prompts for each downstream task. VPT (Jia et al., 2022) is one representative method. One drawback of VPT is that its task-specific prompts are obtained through supervised learning, different from the text prompts used in NLP fields that users give. Another drawback is that it individually searches the prompt length for each task. It is not flexible in applying to a new task. Another direction (Chen et al., 2022b; Houlsby et al., 2019; Jie and Deng, 2022) is to add an adapter alongside the multi-head self-attention (MHSA) or MLP block in ViT (Dosovitskiy et al., 2020). Regardless of whether the structure of the adapter is MLP-like (Houlsby et al., 2019; Chen et al., 2022b) or Convolution-like (Jie and Deng, 2022), these modules overlook leveraging the task-specific information in their efficient module, and the number of trainable parameters is not quite small and offers suboptimal performance.
@@
Though many works are proposed to explore the parameter-efficient tuning techniques, there still lacks the exploration of utilizing the task-specific information in the efficient module design as it is evident that channel bias exists in diverse downstream tasks(Luo et al., 2022). In this paper, we propose a simple and straightforward baseline that aims to find partial channels in the feature map by forwarding the pre-trained model with task-specific images. Such partial tuning enables us only fine-tune a small portion of channels which reduces the tunable parameters heavily. Our experiments surprisingly demonstrate that tuning only a small portion of task-specific channels is sufficient for downstream task adaptation in the low-data regime. As for selecting the task-specific channels, we propose the Class-Aware Importance Score (CAIS) by adopting norm as the criteria to evaluate the channel importance inspired by the classic pruning method (Han et al., 2015). We refer to the selected channels as “salient channels” (SC) since they have higher activation values and are crucial for task performance. This channel selection procedure helps to avoid the training costs by selecting the best structure as in NOAH(Zhang et al., 2022) or the best prompt length as in VPT(Jia et al., 2022). Compared with adapter-based methods (Houlsby et al., 2019; Chen et al., 2022b), selecting partial task-specific channels enables us to achieve lower parameter costs by removing the operations of “downsample” and “nonlinearity” as shown in Fig. 2. We conclude the contributions of our SCT as follows:
@@
We propose a simple baseline with the perspective in partial channel tuning (i.e., salient channel tuning) for addressing the PEFT problem with leveraging task-specific information, achieving promising performance on 19 downstream tasks, domain generalization, and few-shot learning. The experiments demonstrate that tuning a subset of channels in a feature map efficiently for fine-tuning downstream tasks.
It is the first time to reduce the extra parameters into 0.11M level of ViT-B with such competitive performance, which is 780 fewer parameters than the original model.
Our proposed task-specific information storage baseline (SCT) offers a simple solution that requires minimal parameters while preventing the degradation of the original model in downstream domains (full fine-tuning). The low parameter requirements and the ability to adapt with few samples make it a potential task knowledge storage unit for real-world applications.
@@ 2.1 Vision Transformers
Transformer (Vaswani et al., 2017) has demonstrated outstanding results on natural language processing and computer vision tasks. Lots of vision transformers (Chen et al., 2021a; d’Ascoli et al., 2021; Dong et al., 2022; Ali et al., 2021; Fan et al., 2021; Han et al., 2021; Rao et al., 2021; Yuan et al., 2021; Touvron et al., 2021; Liu et al., 2021; Wang et al., 2021; Zhou et al., 2021a) are proposed after the pioneering work ViT (Dosovitskiy et al., 2020). Many of them increase the model size gradually for state-of-the-art results and learn the rich representations by various architectural designs. Noteworthy, most of them are trained on the natural dataset and have the strong potential to be transferred to other domains/tasks. Adopting these models to the downstream tasks alleviates the training difficulty obviously and achieves promising results rapidly.
@@ 2.2 Parameter-Efficient Fine-Tuning Methods
PEFT focuses on adopting a trainable module with a few parameters for fine-tuning. Two lines of PEFT have been proposed recently. On the one hand, applying prompts (Jia et al., 2022; Liu et al., 2022; Xing et al., 2022; Zheng et al., 2022; Nie et al., 2022; Wang et al., 2022; Zhou et al., 2022a, b; Liao et al., 2023; Manli et al., 2022; Zhang et al., 2023b; Zang et al., 2022; Bar et al., 2022) to the backbone networks shows success on several vision tasks. VPT (Jia et al., 2022) first proposes the prompt-based method in computer vision fields. VPT injects the prompts into the transformer layer with a small number of extra parameters. However, one main limitation of VPT is that it relies on hand-crafted selection to determine the optimal prompt length for each task. This can be inflexible when applying the method to new tasks. Another limitation is that the task-specific prompts of VPT are obtained through the training procedure and cannot be given by the user as in the NLP field. In our work, to explicitly reduce the computations of searching the task-specific prompt length, we pass the training images into the backbone network and leverage the task-specific information by determining the salient channels. This procedure only needs one forward propagation without training costs.
@@
On the other hand, adding a residual module (Houlsby et al., 2019; Chen et al., 2022b; Jie and Deng, 2022; Chen et al., 2022a) in the backbone networks also acquires promising results for performance and efficiency. Adapter (Houlsby et al., 2019), a widely used baseline in many tasks (Sung et al., 2022; Pan et al., 2022; Zhang et al., 2023a), proposes an MLP-like module, a successful design that first projects the original dimensional features into a smaller dimension with one nonlinear layer and projects it back to the original dimensions. It vastly reduces the number of parameters. Inspired by its smallest intermediate dimensions, finding a small number of salient channels in a feature map might be enough for the adaptation. Unlike injecting trainable modules into the transformer blocks, LoRA (Hu et al., 2021) optimizes a low-rank decomposition matrix with a low intrinsic dimension to project the query, key, and value matrices used in multiheaded self-attention in ViT. As for NOAH (Zhang et al., 2022), a neural architecture search algorithm incorporates Adapter, LoRA, and VPT into its network search space. NOAH provides a strong baseline for performing consistently well on different datasets. SSF (Lian et al., 2022), a recently proposed strong baseline by only scaling and shifting the features to implement the efficient model tuning. Child-Tuning Xu et al. (2021) updates a subset of parameters of large pre-trained language models via strategically masking out the gradients of the non-child network during the backward process. AdaptFormer Chen et al. (2022b) is an adapter-like method. It explores the efficiency of the adapter-like architecture of vision transformers’ fine-tuning.
@@
Unlike the above methods, our SCT handles the PEFT by leveraging the task-specific information and then proposes the Salient Channel Tuning Module (SCTM) from the new perspective of partial channel tuning, achieving low parameter costs and higher performance than previous methods.
@@ 3.1 Not all channels are equal
It naturally raises a question: Could we find salient channels in a feature map and then only tune these salient channels for efficient tuning? To answer this question, we design a simple class-aware importance score to identify these salient channels and leverage a salient channel tuning module to tune these channels efficiently.
@@ 4 Experiments
This section compares our SCT with other state-of-the-art PEFT baselines on the VTAB-1K benchmark, using ViT and Swin Transformer backbones. In addition, we analyze the channel selection strategy, class-aware importance score, insert position, insert length, and the number of selected channels to verify SCT’s effectiveness further. Beyond evaluating performance on various downstream tasks, we also test SCT’s domain generalization capabilities using four datasets with natural distribution shifts and evaluate its performance in low-data regimes with five datasets in the few-shot scenario.

### v5

@@ Abstract
Pre-trained vision transformers have strong representation benefits to various downstream tasks. Recently, many parameter-efficient fine-tuning (PEFT) methods have been proposed, and their experiments demonstrate that tuning only 1% extra parameters could surpass full fine-tuning in low-data resource scenarios. However, these methods overlook the task-specific information when fine-tuning diverse downstream tasks. In this paper, we propose a simple yet effective method called “Salient Channel Tuning” (SCT) to leverage the task-specific information by forwarding the model with the task images to select partial channels in a feature map that enables us to tune only 1/8 channels leading to significantly lower parameter costs. Experiments on 19 visual transfer learning downstream tasks demonstrate that our SCT outperforms full fine-tuning on 18 out of 19 tasks by adding only 0.11M parameters of the ViT-B, which is 780 fewer than its full fine-tuning counterpart. Furthermore, experiments on domain generalization and few-shot classification further demonstrate the effectiveness and generic of our approach. The code is available at https://github.com/showlab/SCT
@@ 1 Introduction
Large vision transformers (ViT) have achieved remarkable success in computer vision tasks (Dosovitskiy et al., 2020; Liu et al., 2021; Yuan et al., 2022; Zhou et al., 2021a; Carion et al., 2020; Li et al., 2021; Strudel et al., 2021) by using large-scale training data, such as ImageNet21K and JFT-300M, to create strong representations. However, downstream recognition tasks often lack sufficient data to train from scratch. Therefore, transferring knowledge from pre-trained ViT models can significantly reduce training difficulty and produce promising results. End-to-end full fine-tuning is a widely used way to inherit these robust representations, but it faces two challenges. First, large models are prone to overfitting when tuning their massive weights on small downstream training data. Second, ViT models are too large to store all weights for each downstream task, making it infeasible to deploy fine-tuned models on resource-limited devices.
@@
To mitigate the above two challenges, some works propose to tune a subset of parameters (zaken2021bitfit) or adopt an external trainable module (Houlsby et al., 2019; Hu et al., 2021) to preserve the knowledge learned from pre-trained models. As for tuning a subset of parameters, there are two representative approaches: tuning the classification head (Mahajan et al., 2018; Jia et al., 2021; Chen et al., 2021b) and tuning bias term (Cai et al., 2020). Tuning the classification head involves freezing the weights of the backbone network and updating only the linear head, whereas tuning the bias term entails unfreezing the bias term within the backbone network. Both approaches result in inferior performance.
Recently, some studies proposed to address a new task named Parameter-Efficient Fine-Tuning (PEFT) by leveraging an external trainable module for model adaptation. One type of method includes integrating learnable prompts into the model input for each downstream task, exemplified by the pioneering work, VPT(Jia et al., 2022). However, VPT encounters several challenges. First, its task-specific prompts are derived through supervised learning, which differs from the user-provided text prompts commonly used in the NLP field. Second, VPT requires searching for the best prompt length for each task which lacks flexibility and is impractical when encountering new downstream tasks. Another type of work is to add an adapter (Chen et al., 2022b; Houlsby et al., 2019; Jie and Deng, 2022) alongside the multi-head self-attention (MHSA) or MLP block and address the whole features equally without considering the task-specific information into the efficient module design. Moreover, the number of trainable parameters is not quite small.
@@
By investigating these approaches, utilizing the task-specific information in the efficient module design is still underexplored as it is highlighted in a recent study (Luo et al., 2022) that channel bias exists in diverse downstream tasks. In this paper, to address the PEFT problem, we propose a simple baseline to involve task-specific information by forwarding the pre-trained model with downstream images. Furthermore, we proceed to select a small portion of channels to achieve efficient fine-tuning, significantly reducing the parameter costs. Our experiments in Sec.4.2 reconfirm that tuning only a small portion of task-specific channels is sufficient for downstream task adaptation in the low-data regime. To achieve the goal of obtaining these task-specific channels, we propose a Class-Aware Importance Score (CAIS) by adopting norm as the criteria to evaluate the channel importance inspired by the classic pruning method (Han et al., 2015). We refer to the selected channels as “salient channels” (SC) since they have higher activation values and are crucial for task performance. It is noted that channel selection can avoid the training costs by selecting the best structure as in NOAH(Zhang et al., 2022) or the best prompt length as in VPT(Jia et al., 2022). Moreover, compared with adapter-based methods (Houlsby et al., 2019; Chen et al., 2022b), selecting partial task-specific channels enables us to achieve lower parameter costs by removing the operations of “downsample” and “nonlinearity” as shown in Fig. 2.
In summary, the contributions are summarized as follows:
@@
We propose a simple baseline with a new perspective in partial channel tuning (i.e., salient channel tuning) for addressing the PEFT task to leverage task-specific information, achieving promising performance on 19 visual transfer learning tasks, 4 domain generalization tasks, and 5 few-shot learning tasks. The experiments demonstrate that tuning a subset of channels in a feature map is efficient and generic.
To the best of our knowledge, it is the first time to reduce the parameter costs to 0.11M (780 fewer parameters than the full model size) level of ViT-B while achieving the best average performance compared with other methods.
@@ 2.1 Vision Transformers
Transformer (Vaswani et al., 2017) has demonstrated outstanding results on natural language processing and computer vision tasks. Lots of vision transformers (Chen et al., 2021a; d’Ascoli et al., 2021; Dong et al., 2022; Ali et al., 2021; Fan et al., 2021; Han et al., 2021; Rao et al., 2021; Yuan et al., 2021; Touvron et al., 2021; Liu et al., 2021; Wang et al., 2021; Zhou et al., 2021a) are proposed after the pioneering work ViT (Dosovitskiy et al., 2020). Many of them increase the model size gradually for state-of-the-art results and learn the rich representations by various architectural designs. It is noted that most of them are trained on the natural dataset and have the strong potential to be transferred to other domains/tasks. Moreover, adopting these models to the downstream tasks is able to alleviate the training difficulty and achieve promising results.
@@ 2.2 Parameter-Efficient Fine-Tuning Methods
PEFT focuses on adopting a trainable module with a few parameters for fine-tuning. Two lines of PEFT have been proposed recently. On the one hand, applying prompts (Jia et al., 2022; Liu et al., 2022; Xing et al., 2022; Zheng et al., 2022; Nie et al., 2022; Wang et al., 2022; Zhou et al., 2022a, b; Liao et al., 2023; Manli et al., 2022; Zhang et al., 2023b; Zang et al., 2022; Bar et al., 2022) to the backbone networks shows success on several vision tasks. VPT (Jia et al., 2022) first proposes the prompt-based method in computer vision fields by injecting the prompts into each transformer layer. However, one main limitation of VPT is that it relies on hand-crafted selection to determine the optimal prompt length for each task. This is inflexible when applying to a new downstream task. Another limitation is that the task-specific prompts of VPT are obtained through the training procedure and cannot be given by the user as in the NLP field. In our work, to explicitly reduce the computations of searching the task-specific prompt length, we pass the training images into the backbone network and leverage the task-specific information by determining the salient channels. This procedure only needs one forward propagation without training costs.
@@
On the other hand, adding a residual module (Houlsby et al., 2019; Chen et al., 2022b; Jie and Deng, 2022; Chen et al., 2022a) in the backbone networks also acquires promising results for performance and efficiency. Adapter (Houlsby et al., 2019), a widely used baseline in many tasks (Sung et al., 2022; Pan et al., 2022; Zhang et al., 2023a), proposes an MLP-like module, a successful design that first projects the original dimensional features into a smaller dimension with one nonlinear layer and projects it back to the original dimensions. It vastly reduces the number of parameters. Inspired by its smallest intermediate dimensions, finding a small number of salient channels in a feature map might be enough for the adaptation. Unlike injecting trainable modules into the transformer blocks, LoRA (Hu et al., 2021) optimizes a low-rank decomposition matrix with a low intrinsic dimension to project the query, key features. As for NOAH (Zhang et al., 2022), a neural architecture search algorithm incorporates Adapter, LoRA, and VPT into its network search space. NOAH provides a strong baseline for performing consistently well on different datasets. SSF (Lian et al., 2022), a recently proposed strong baseline by only scaling and shifting the features to implement the efficient model tuning. Child-Tuning Xu et al. (2021) updates a subset of parameters of large pre-trained language models via strategically masking out the gradients of the non-child network during the backward process. AdaptFormer Chen et al. (2022b) is an adapter-like method. It explores the efficiency of the adapter-like architecture of vision transformers’ fine-tuning.
@@
Unlike the above methods, our SCT handles the PEFT by leveraging the task-specific information into the method design and then proposes the Salient Channel Tuning, achieving lower parameter costs and higher performance than previous methods.
@@ 3.1 Not all channels are equal
It naturally raises a question: Could we find salient channels in a feature map and then only tune these salient channels for efficient tuning? To answer this question, we design a simple class-aware importance score to identify these salient channels and leverage a salient channel tuning module for efficient fine-tuning.
@@ 4 Experiments
This section compares our SCT with other state-of-the-art PEFT baselines on the VTAB-1K benchmark, using ViT and Swin Transformer backbones. In addition, we analyze the channel selection strategy, class-aware importance score, insert position, insert length, and the number of selected channels to verify SCT’s effectiveness further. Beyond evaluating performance on visual transfer learning tasks, we also validate the capabilities of four domain generalization datasets and test the performance of SCT in the low-data scenarios on five datasets in the few-shot setting.